OpenStudy (anonymous):

A 0.150 kg air-track glider moving at 1.75 m/s bumps into a 0.350 kg glider at rest. Part 1: Find the total kinetic energy after collision if the collision is elastic. Part 2: Find the total kinetic energy after collision if the collision is completely inelastic. ***How can I solve this and show the work for it? :/

OpenStudy (johnweldon1993):

Okay hmm, this is taking me back to physics 1...argh 3 year old memory haha well basically since it would be elastic (part a) there would be no loss of kinetic energy..so we have \[\large K.E_{initial} = K.E_{final}\] so what the the total kinetic energy in the beginning? \[\large K.E_{air glider 1} + K.E_{air glider 2}\] \[\large \frac{1}{2}mv^2 + \frac{1}{2}mv^2\]...

OpenStudy (anonymous):

hahaha :P and oaky! so from there, we know it is elastic, so we have to find the total kinetic energy? what would we plug in? :O

OpenStudy (johnweldon1993):

It is elastic so we know whatever we get from the initial total kinetic energy is our answer because it is the same as AFTER the collision So before we have For air glider 1 \[\large K.E = \frac{1}{2}mv^2 = \frac{1}{2}(0.150kg)(1.75m/s)^2 = ?\] And for air glider 2 \[\large K.E = \frac{1}{2}mv^2 = \frac{1}{2}(0.350kg)(0m/s)^2 = 0\] Right? and the total kinetic energy is the sum of those...so what is \[\large \frac{1}{2}(0.150kg)(1.75m/s)^2 + 0 = ?\]

OpenStudy (anonymous):

ahh okay!! so it would equal 0.13125? not sure if i calculated that correctly?

OpenStudy (johnweldon1993):

Argh you're gonna make me calculate it? >.< haha no jk hang on...

OpenStudy (johnweldon1993):

Nope not quite what I get...

OpenStudy (johnweldon1993):

You didn't square the velocity...

OpenStudy (anonymous):

haha im not sure... a lot of times i tend to do something wrong lol like i forget to change something in my calc :/ i get confused ohhh darn :/ okay, let me try again :)

OpenStudy (anonymous):

ohh so it would be 0.2296875? :/

OpenStudy (johnweldon1993):

Indeed....but I think just rounding to 0.23 is fine

OpenStudy (anonymous):

ahh okay yay!! :) so that is the total kinetic energy for part 1? 0.23? also, i forgot, what are the units again? :/

OpenStudy (johnweldon1993):

Well here we have kg * (m/s)^2 would be kg m^2 / s^2 *kilogram- meters squared per seconds squared

OpenStudy (johnweldon1993):

And a hint in case you hate all that 1 joule ( J ) is equal to 1 kg m^2 / s^2 So just 0.23 J would be good

OpenStudy (anonymous):

ohh okay! :) awesome!! i will try to remember that :P so onto part 2? :O

OpenStudy (johnweldon1993):

Sure....so part 2 we have an INELASTIC collision...meaning the 2 objects stick together after they collide right?

OpenStudy (anonymous):

yes! :)

OpenStudy (johnweldon1993):

So...we start off with the normal kinetic energy and also the original momentum 1/2 mv^2 and mv then they collide (boom!) lol now they are stuck together

OpenStudy (anonymous):

ok:) haha boom, got it! :P

OpenStudy (johnweldon1993):

Thanks to the conservation of momentum...the momentum before and after the collision are the same... \[\large m_1 v_1 = (m_1 + m_2) v_2 \] why is it (m1 + m2) over on the right? because they are now stuck together...so the total mass would be the mass of both good so far?

OpenStudy (anonymous):

ahh okay:) yes, good :)

OpenStudy (johnweldon1993):

And the kinetic energy at the end is going to be basically the same \[\large K.E = \frac{1}{2}(m_1 + m_2)v_2^2 \] Since this is AFTER the collision...we use the final velocity (v2) and we still have the 2 masses because they are stuck together....still good?

OpenStudy (anonymous):

yes:)

OpenStudy (johnweldon1993):

So we have 2 equations The momentum: \(\large m_1 v_1 = (m_1 + m_2)v_2\) and the kinetic energy \(\large \frac{1}{2}(m_1 + m_2)v_2 ^2\) So what we need, is to know the final velocity....looking at the momentum equation...how would we solve for the v2?

OpenStudy (anonymous):

isolate it?

OpenStudy (johnweldon1993):

Right...so we would get \[\large v_2 = \frac{m_1 v_1}{m_1 + m_2}\] right?

OpenStudy (anonymous):

right :)

OpenStudy (johnweldon1993):

Now let's just plug that into our kinetic energy equation and we will finally be ready to solve \[\large K.E = \frac{1}{2}(m_1 + m_2)(\frac{m_1v_1}{m_1 + m_2})^2\] man that's a long one...but we have everything we need...we know both the masses...and we know the original velocity...so just plug everything in and we will have the answer

OpenStudy (johnweldon1993):

and yes I'll do it too to check :P

OpenStudy (anonymous):

haha okie :) so we will get this? hopefully i am plugging in correctly lol KE=1/2(0.150+0.350)((0.150*1.75)/(0.150+0.350))^2 ? and that equals 7686.905625 ? :/ ooh something looks weird :( not sure what i did weirdly and incorrectly though :/

OpenStudy (johnweldon1993):

Oh boy no lol You did indeed plug everything in right though \[\large \frac{1}{2}(0.150 + 0.350)(\frac{0.150\times 1.75}{0.150 + 0.350})^2\] \[\large \frac{1}{2}(0.5)(\frac{0.2625}{0.5})^2\] \[\large \frac{1}{2}(0.5)(0.275625)\] \[\large 0.0689\]

OpenStudy (anonymous):

oh yay at least i plugged in right! haha oops i was really off :( i'm not very good with the calculator!! :( so it would be 0.0689 J ? that is the final total kinetic energy for part b? :O

OpenStudy (johnweldon1993):

Correct :)

OpenStudy (anonymous):

yay!! thank you so much! :)

OpenStudy (johnweldon1993):

Not a problem...and now...I must bid you farewell D: lol

OpenStudy (anonymous):

:)